OpenStudy (anonymous):

Solve the system, using substitution. What is the x coordinate of the solution? x = 2y - 2 and 3x - y = 14 A. 6 B. 4 C. 1 D. 0 E. no solution

OpenStudy (bibby):

replace x with (2y-2) in the 2nd equation

OpenStudy (anonymous):

so the answer would be a?

OpenStudy (bibby):

\(x = 2y - 2 \\ 3x - y = 14\implies 3(2y-2) - y = 14 \) solve for y, plug it into one of the equations to find x

OpenStudy (anonymous):

4?

OpenStudy (bibby):

yeah, A seems right

OpenStudy (anonymous):

A?? not B?

OpenStudy (bibby):

\(3(2y−2)−y=14\\6y-6-y=14\\6y-y-6=14\\5y-6=14\\5y=20\\y=4\) then we plug it back into one of the original equations

OpenStudy (bibby):

god, I hate this site

OpenStudy (anonymous):

im lost

OpenStudy (bibby):

so we have \(x = 2y - 2 \\y=4\) \(x=2y-2\\x=2(4)-2\)

OpenStudy (bibby):

do you know how to solve by substitution?

OpenStudy (anonymous):

no

OpenStudy (bibby):

there's a solution (x,y) to this system ``` x = 2y - 2 3x - y = 14 ``` solving by substitution involves replacing one of the variables in one equation with part of another equation. basically, we know that x=2y-2 using that we can plug it into the second equation and solve for the y coordinate of the system's solution then we plug the number we get for the y into that same equation and x=8-2=6

OpenStudy (anonymous):

Why didnt you just tell me i was right when i wrote A...

OpenStudy (bibby):

because you could have guessed it

OpenStudy (bibby):

I'm sorry for putting in 5 minutes of my time, free of charge, to try and teach you something

OpenStudy (bibby):

Learn to be respectful and grateful

OpenStudy (anonymous):

¯\_(ツ)_/¯

OpenStudy (bibby):

http://www.alexdantas.net/lenny/